# OE_^ OmOE_^ OmOE_^ Omm!OE_^ mmmmOE_^

### v1

@@ 6.3. Hyperbolic fractal uncertainty principle
$\Theta:\tilde{U}\times\left[\tfrac{1}{4}-\tfrac{2\nu_{1}}{C_{1}^{2}},4+\tfrac{2\nu_{1}}{C_{1}^{2}}\right]\rightarrow W^{\prime\prime},\quad(u,v,\lambda)\mapsto\kappa_{0}^{-}(f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{+}v}e^{U^{-}u}(\tilde{x},\tilde{\xi},\tilde{\Xi})))),$
@@
Let $\{M_{k}\}$ be a finite covering of $\mathbb{S}^{n}$ by balls of fixed radius $1/2$, each centered at $y_{k}\in\mathbb{S}^{n}$. We view $\mathbb{S}^{n}$ as a subset of $\mathbb{R}^{n+1}$ and identify each hyperplane tangent to $y_{k}$ with $\mathbb{R}^{n}$. Then let $\psi_{k}:M_{k}\rightarrow\mathbb{R}^{n}$ be gnomonic projection onto the hyperplane tangent to $y_{k}$. More specifically, for $y\in M_{k}$, $\psi_{k}(y)$ is the intersection of the the hyperplane tangent to $y_{k}$ and the line going through the center of $\mathbb{S}^{n}$ and $y$. We know $\psi_{k}(M_{k})=X$, where $X\subset\mathbb{R}^{n}$ is a ball. There exists some $C_{2}>0$ such that for $x,y\in X$ and all $k$,
@@
We begin by showing the ball porosity of $\psi_{k}(\Omega_{+}\cap M_{k})$. Let $R\in[h^{\varrho},1]$. We examine $B_{R/2}(r_{0})$ for some $r_{0}\in\mathbb{R}^{n}$. Either $B_{R/2}(r_{0})$ contains a ball of radius $R/4$ contained inside $X$ or $B_{R/2}(r_{0})$ contains a ball of radius $R/4$ that does not intersect $X$. Since $\nu/2C_{2}\leq 1/4$, it suffices to examine $B_{R/4}(r_{0})$ contained in $X$.
@@
$B_{\nu R/2C_{2}}(\psi_{k}(y_{1}))\cap X\subset\psi_{k}\left(B_{\nu R/2C_{2}}(y_{1})\cap M_{k}\right).$
@@
We now show the line porosity of $\psi_{k}(\Omega_{-}\cap M_{k})$. Let $I$ be a line segment of length $R\in[h^{\varrho},1]$. Either $I\cap X$ contains a line segment of length $R/2$ or $I\setminus X$ contains a line segment of length $R/4$. Since $\nu/2C_{2}\leq 1/4$, it suffices to assume $I$ is a line segment of length $R/2$, contained in $X$.
By  (6.32), $\psi_{k}^{-1}(I)$ is a segment of a geodesic on $\mathbb{S}^{n}$ of length at least $R/2C_{2}$. Recall $C_{1},\nu_{1},K_{1}$ from Lemma 6.12. For $h$ sufficiently small, $C_{1}K_{1}h^{\rho}/\nu_{1}<h^{\varrho}/2C_{2}\leq R/2C_{2}\leq 1$. Therefore by Lemma  6.12, there exists some $y_{1}\in\psi_{k}^{-1}(I)$ such that $B_{\nu R/2C_{2}}(y_{1})\cap\Omega_{-}=\emptyset$. Clearly, $\psi_{k}(y_{1})\in I$. From  (6.32), we know $B_{\nu R/2C_{2}}(\psi_{k}(y_{1}))\cap X\subset\psi_{k}(B_{\nu R/2C_{2}}(y_{1})\cap M_{k})$. Therefore, $B_{\nu R/2C_{2}}(\psi_{k}(y_{1}))\cap\psi_{k}(\Omega_{-}\cap M_{k})=\emptyset$. We conclude that $\psi_{k}(\Omega_{-}\cap M_{k})$ is $\nu/2C_{2}$-porous on lines from scales $h^{\varrho}$ to $1$. ∎

### v2

@@ 6.3. Hyperbolic fractal uncertainty principle
$\Theta:\tilde{U}\times\left[\tfrac{1}{4}-\tfrac{2\nu_{1}}{C_{1}^{2}},4+\tfrac{2\nu_{1}}{C_{1}^{2}}\right]\rightarrow W^{\prime\prime},\quad(u,v,\lambda)\mapsto\kappa_{0}^{-}(f_{\lambda}(\pi_{S}(e^{\mathcal{V}^{+}v}e^{\mathcal{U}^{-}u}(\tilde{x},\tilde{\xi},\tilde{\Xi})))),$
@@
Let $\{M_{k}\}$ be a finite covering of $\mathbb{S}^{n}$ by balls of fixed radius $1/2$, each centered at $y_{k}\in\mathbb{S}^{n}$. We view $\mathbb{S}^{n}$ as a subset of $\mathbb{R}^{n+1}$ and identify each hyperplane tangent to $y_{k}$ with $\mathbb{R}^{n}$. Then let $\psi_{k}:M_{k}\rightarrow\mathbb{R}^{n}$ be gnomonic projection onto the hyperplane tangent to $y_{k}$. More specifically, for $y\in M_{k}$, $\psi_{k}(y)$ is the intersection of the the hyperplane tangent to $y_{k}$ and the line going through the center of $\mathbb{S}^{n}$ and $y$. We know
$\psi_{k}(M_{k})=B,$
where $B\subset\mathbb{R}^{n}$ is a ball. There exists some $C_{2}>0$ such that for $x,y\in B$ and all $k$,
@@
We begin by showing the ball porosity of $\psi_{k}(\Omega_{+}\cap M_{k})$. Let $R\in[h^{\varrho},1]$. We examine $B_{R/2}(r_{0})$ for some $r_{0}\in\mathbb{R}^{n}$. Either $B_{R/2}(r_{0})$ contains a ball of radius $R/4$ contained inside $B$ or $B_{R/2}(r_{0})$ contains a ball of radius $R/4$ that does not intersect $B$. Since $\nu/2C_{2}\leq 1/4$, it suffices to examine $B_{R/4}(r_{0})$ contained in $B$.
@@
$B_{\nu R/2C_{2}}(\psi_{k}(y_{1}))\cap B\subset\psi_{k}\left(B_{\nu R/2C_{2}}(y_{1})\cap M_{k}\right).$
@@
We now show the line porosity of $\psi_{k}(\Omega_{-}\cap M_{k})$. Let $I$ be a line segment of length $R\in[h^{\varrho},1]$. Either $I\cap B$ contains a line segment of length $R/2$ or $I\setminus B$ contains a line segment of length $R/4$. Since $\nu/2C_{2}\leq 1/4$, it suffices to assume $I$ is a line segment of length $R/2$, contained in $B$.
By  (6.32), $\psi_{k}^{-1}(I)$ is a segment of a geodesic on $\mathbb{S}^{n}$ of length at least $R/2C_{2}$. Recall $C_{1},\nu_{1},K_{1}$ from Lemma 6.12. For $h$ sufficiently small, $C_{1}K_{1}h^{\rho}/\nu_{1}<h^{\varrho}/2C_{2}\leq R/2C_{2}\leq 1$. Therefore by Lemma  6.12, there exists some $y_{1}\in\psi_{k}^{-1}(I)$ such that $B_{\nu R/2C_{2}}(y_{1})\cap\Omega_{-}=\emptyset$. Clearly, $\psi_{k}(y_{1})\in I$. From  (6.32), we know $B_{\nu R/2C_{2}}(\psi_{k}(y_{1}))\cap B\subset\psi_{k}(B_{\nu R/2C_{2}}(y_{1})\cap M_{k})$. Therefore, $B_{\nu R/2C_{2}}(\psi_{k}(y_{1}))\cap\psi_{k}(\Omega_{-}\cap M_{k})=\emptyset$. We conclude that $\psi_{k}(\Omega_{-}\cap M_{k})$ is $\nu/2C_{2}$-porous on lines from scales $h^{\varrho}$ to $1$. ∎